OpenStudy (linda):

algebraically and then graphically find all zeroes of the polynomial function x^4-5x^2+4=0

OpenStudy (anonymous):

\[(x^2-4)(x^2-1)=0\] \[(x^2-4)=0\] \[x=\pm2\] \[x^2-1=0\] \[x=\pm1\]